OpenStudy (anonymous):

What is the distance between the two points (5, –3) and (3, 9)? SquareRoot 148 SquareRoot 165 SquareRoot 181 Squareroot 198

Nnesha (nnesha):

what's the formula to find distance ?

OpenStudy (anonymous):

@Nnesha I don't know, that's my issue o-o

Nnesha (nnesha):

ahh alright \[\huge\rm \sqrt{(x_2 -x_1)^2 + (y_2-y_1)^2}\] okay here is a formula where x and y values are \[\huge\rm (x_1 ,y_1) (x_2 ,y_2)\] plug in x and y values

OpenStudy (anonymous):

Okay so x2 - x1 = 3 - 5 = -2 ys - y1 = 9 - (-3) = 12 12 +-2 = 10 ??? Did I do that right?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

yep -2 and 12 now you can't add them yet bec it's (x_2 - x_1)^2 <square so first take square of both number then add them

Nnesha (nnesha):

\[\huge\rm \sqrt{(x_2 -x_1)^\color{ReD}{2} + (y_2-y_1)^\color{red}{2}}\] \[\huge\rm \sqrt{(-2)^\color{ReD}{2} + (12)^\color{red}{2}}\]

OpenStudy (anonymous):

^2 of -2 = 4 ^2 of 12 = 144 So add 4 + 144 ? = 148

Nnesha (nnesha):

yes right sqrt of 148

OpenStudy (anonymous):

Okay thank you :D

Nnesha (nnesha):

my pleasure :-)

Nnesha (nnesha):

btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{gold}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)

OpenStudy (anonymous):

Thank you :D